OpenStudy (anonymous):

A high-altitude spherical weather balloon expands as it rises due to the drop in atmospheric pressure. Suppose that the radius r increases at the rate of 0.16 inches per second and that r = 38 inches at time t = 0. Determine the equation that models the volume V of the balloon at time t and find the volume when t = 280 seconds. actually the last one this time. id really appreciate it!

OpenStudy (kesumonu):

let r(t)=0.16*t+38 now V=(4/3)(pi)(r(t))^3

OpenStudy (kesumonu):

find volume at t=280 by substituting r(280)=.16*280 38=82.8

OpenStudy (anonymous):

V(t) = ; 2,377,823.53 in.3 or V(t) = ; 1,129,910.14 in.3

OpenStudy (kesumonu):

that right

OpenStudy (anonymous):

which one??

OpenStudy (kesumonu):

1st one

OpenStudy (anonymous):

YIPPEEEE thanks a bunch!

OpenStudy (kesumonu):

u study at college?